OpenStudy (anonymous):

If 1000 square centimeters of material is available to make a box with a square base and an open top, find the largest possible volume of the box.

OpenStudy (mertsj):

The area of the base is x^2> The height is h. Each side of the box has area xh. There are 4 sides of the box so the total surface area of the box is x^2+4xh and that is equal to 1000. Solve that equation for h: x^2+4xh = 1000 h = (1000-x^2)/4x so the Volume = x^2[(1000-x^2)/4x] Simplify and get V = 250x-x^3/4 The volume will be a maximum when its first derivative is 0. V' = 250-3/4x^2 Set to 0 and solve. x=18.26 Now plug into the volume function to find the maximum volume: V=250(18.26)-(18.26)^3/4 V= 4564.35 - 1522.10 =3042.25

OpenStudy (anonymous):

Thank you so much!!! Can you by any chance solve this problem as well? I would reallyyy appreciate it. A cylinder is inscribed in a right circular cone of height 3.5 and radius (at the base) equal to 5.5. What are the dimensions of such a cylinder which has maximum volume?

OpenStudy (mertsj):

I will try as soon as I get done watching Hannity's interview with Newt

OpenStudy (anonymous):

Ok my assignment is due in less than an hour though. Again thanks a lotttt

OpenStudy (mertsj):

5.5/3.5=r/(3.5-x) Solve for r r=(38.5-11x)/7 Plug into V formula v = pir^2h V= pi[(38.5-11x^2/7]^2(x) =pix/49(1482.25x-847x^2+121x^3) V' = pi/49(1482.25-1694x+363x^2) Set to 0 and use the quadratic formula

OpenStudy (anonymous):

I only need the final answer. No explanation needed.

OpenStudy (anonymous):

Please and thank you so much. Really appreciate it.

OpenStudy (mertsj):

you will get (1694+ - 847)/726 V = 15.7

OpenStudy (anonymous):

I need the dimensions not the volume. Like what is the height and radius when the volume is maximized???

OpenStudy (mertsj):

r=3.66

OpenStudy (anonymous):

not correct :(

OpenStudy (anonymous):

due in 5 mins pls help

OpenStudy (mertsj):

h=2.33

OpenStudy (anonymous):

both incorrect...

OpenStudy (mertsj):

How do you know?

OpenStudy (anonymous):

i submit it online and it tells me if the answer is incorrect

OpenStudy (anonymous):

but i can try again

OpenStudy (mertsj):

Should be close

OpenStudy (anonymous):

A baseball team plays in he stadium that holds 68000 spectators. With the ticket price at 9 the average attendence has been 29000. When the price dropped to 8, the averege attendence rose to 34000. a) Find the demand function p(x), where x is the number of the spectators. (assume p(x) is linear). Im really sorry im annoying you with all these questions but im stressed because its due now...

OpenStudy (mertsj):

let me know if you find out the right answer.

OpenStudy (anonymous):

ok do you know what the demand function is?

OpenStudy (mertsj):

p(x)=5000x-16000 ???

OpenStudy (mertsj):

p(x)=-5000x + 64000 I think the slope is negative

OpenStudy (mertsj):

p(x) = -5000x +74000